material Assigment × G what does roundngtoone dec × 1 + Upgrade to macOS Mojave 734 m/courses/14882 Get Dark Made, Stacka, new apps a new Mac App Store a) Create a new two-way table with row and column totals added HS Some College Total 176 196 789 1812 557 |[20 ] [26 ] Dont know 32 78 Total eTextbook and Media Your answer is partlially correct Ib) Find the percent who agree that there is only one true love for each education level. Round your answers to one decimal place. HS College Percent % Does there seem to be an assoclation between education level and agreement with the statement? 名。 o0
Answer #1

For HS = 363/940 * 100 = 38.617%

For some = 176/668 * 100 = 26.347%

For college = 196/1017 * 100= 19.272 %